# Zhuo Zhang College of Computer Science and Technology

### v2

@@ 1.3.1 Proper Orthogonal Decomposition and Galerkin Projection
The Proper Orthogonal Decomposition (POD), also known as Principal Component Analysis or Karhunen-Loève expansion, forms the cornerstone of traditional ROM approaches (Holmes1996; Berkooz1993). Given snapshot solutions $\{u(\cdot,\cdot;\mu_{i})\}_{i=1}^{N}$ collected at training parameters, POD constructs an optimal low-dimensional approximation subspace.
@@ 1.3.2 Reduced Basis Methods
Reduced basis (RB) methods (Quarteroni2015; Hesthaven2016; Rozza2008) extend POD through sophisticated greedy algorithms for parameter space exploration. The key innovation is adaptive selection of snapshot parameters to maximize approximation quality across $\mathcal{P}$.
@@ 1.3.3 Sparse Grid and Multi-Level Methods
For higher-dimensional parameter spaces, sparse grid methods based on Smolyak’s algorithm (Bungartz2004) provide a middle ground between full tensor-product grids and random sampling. The key idea is anisotropic refinement, placing more collocation points along directions of high solution variability.
Multi-level Monte Carlo (MLMC) (Giles2015) tackles high-dimensional uncertainty quantification by combining low-fidelity (coarse mesh) solutions with high-fidelity corrections, achieving variance reduction without prohibitive cost.
@@ 2.1.1 Foundational Framework
Physics-informed neural networks, pioneered by Raissi et al. (Raissi2019) in their seminal 2019 work, represent a breakthrough in integrating physical laws with data-driven learning. The core innovation lies in encoding PDE residuals directly into the loss function, enabling networks to discover solutions satisfying both data constraints and governing equations.
@@ 1. Inverse Problems and Parameter Identification
The pioneering work of Raissi et al. (Raissi2019), building on foundational deep learning approaches for PDEs (Han2018; Sirignano2018; Berg2018; Weinan2019), demonstrated PINNs’ remarkable capability for inverse problems: inferring unknown parameters from sparse observational data. Complementary methodological developments (Raissi2018; Cuomo2022; Lu2022) have further established PINNs as powerful tools for parameter identification. This capability is particularly valuable when parameters represent hidden physical properties.
@@
Case Study: Hidden Fluid Mechanics (Raissi2020Science): Raissi et al. demonstrated parameter identification for the Navier-Stokes equations from scattered velocity measurements. Given noisy observations of the velocity field, the method simultaneously reconstructs the full flow field and infers the Reynolds number, achieving relative errors below 1% with measurements at only 0.1% of the domain points.
@@ 2. Adaptive Sampling Strategies
Residual-Based Adaptive Refinement (RAR): At each training iteration, compute residuals at all collocation points and preferentially sample from high-residual regions (Daw2022; Gao2023; Lu2021Adaptive; Wu2020):
@@
Recent work by Wu et al. (Wu2023) demonstrated 50-70% reductions in required training data through parameter space active learning (Lookman2023; Wu2020) for Reynolds number-dependent flows.
Advanced training methods have improved PINN performance. Wang et al. (Wang2025Gradient) introduced gradient alignment from a second-order optimization perspective, dynamically adjusting loss weights based on gradient alignment to achieve 30-50% training time reductions. De Ryck et al. (DeRyck2024Operator) provided an operator preconditioning perspective on physics-informed training. Hwang and Lim (Hwang2024DualCone) developed dual cone gradient descent for handling conflicting gradients.
Hao et al. (Hao2024PINNacle) released PINNacle at NeurIPS 2024, providing standardized PINN evaluation across 15 canonical PDEs with 20+ method variants. Miyagawa and Yokota (Miyagawa2024Functional) extended PINN theory to functional differential equations with convergence guarantees.
@@ 3. Multi-Task and Transfer Learning
Desai et al. (Desai2021), along with comparative studies (Lu2022; Geneva2022), showed that transfer learning reduces training iterations by up to 80% when adapting PINNs across Reynolds numbers.
@@ 2.1.3 Recent Advances in PINN Architecture and Training
Cao2025IllConditioning established a fundamental connection between PINN ill-conditioning and the Jacobian matrix condition number of the PDE system. By constructing controlled systems that adjust the condition number while preserving solutions, they demonstrated that reducing the Jacobian condition number leads to faster convergence and higher accuracy. This breakthrough enabled the first successful PINN simulation of three-dimensional flow around the M6 wing at Reynolds number 5,000, representing a significant milestone for industrial-complexity problems. Zhang2025P2PINN built upon their foundation, applying it to the parametric domain.
@@
To mitigate the curse of dimensionality (CoD) in multi-dimensional PDEs, several separable architectures have been proposed. These methods primarily focus on decomposing the high-dimensional problem. Cho2023Separable, for instance, introduced Separable Physics-Informed Neural Networks (SPINN), which operates on a per-axis basis to dramatically reduce network forward passes and memory overhead, enabling training with over $10^{7}$ collocation points. Similarly, zhang2025dfs proposed the Dynamic Feature Separation PINN (DFS-PINN), which employs an innovative input-decoupling and dynamic interaction mechanism to achieve significant computational savings. While these methods leverage separation to tackle dimensionality, the concept has also been proven effective against spectral bias. xiong32025separated proposed the Separated-Variable Spectral Neural Network (SV-SNN), which decomposes multivariate functions into products of univariate functions integrated with adaptive Fourier features, specifically to capture high-frequency components effectively.
@@
A paradigm shift in neural architecture comes from Kolmogorov-Arnold Networks (KAN), which replace fixed activation functions with learnable spline-based functions. Wang2024KINN proposed Kolmogorov-Arnold-Informed Neural Networks (KINN), systematically comparing KAN and MLP across various PDE formulations including strong form, energy form, and inverse form. KINN demonstrates significant advantages in accuracy and convergence for multi-scale, singularity, stress concentration, and nonlinear hyperelasticity problems, offering better interpretability with fewer parameters.
Building on this foundation, Jacob2024SPIKANs introduced Separable Physics-Informed Kolmogorov-Arnold Networks (SPIKANs), applying the separation of variables principle to KANs. By decomposing problems such that each dimension is handled by an individual KAN, SPIKANs drastically reduce computational complexity without sacrificing accuracy, particularly effective for higher-dimensional PDEs where collocation points grow exponentially with dimensionality.
Most recently, Zhang2025LegendKINN proposed Legend-KINN, integrating Legendre orthogonal polynomials into the KAN architecture combined with pseudo-time stepping in backpropagation. Legend-KINN achieves 1–3 orders of magnitude faster convergence than both KAN and MLP under identical parameter settings. Further exploring this direction, xiong22025j developed J-PIKAN, a framework based on Jacobi orthogonal polynomials.
@@ 2.1.4 Advantages and Limitations: A Critical Analysis
4. Flexible Uncertainty Quantification: Bayesian extensions like B-PINNs (Yang2021) and related uncertainty quantification approaches (Yang2020Bayesian; Psaros2023; Lakshminarayanan2017; Tripathy2018) naturally quantify parameter and prediction uncertainties without requiring multiple forward solves.
@@
2. Training Instability: Balancing multiple loss terms in Eq. equation 13 remains challenging. Stiff PDEs, multi-scale problems, and high Reynolds number flows often exhibit training instabilities, with loss components competing rather than cooperating (Krishnapriyan2021; Wang2022).
@@
5. Spectral Bias: Neural networks exhibit spectral bias toward learning low-frequency components of solutions (Rahaman2019). For parametric problems with high-frequency features (shocks, boundary layers) that vary with parameters, this bias necessitates very deep networks or specialized architectures.
@@ 2.2.1 Mathematical Foundations of Operator Learning
Chen and Chen (Chen1995) established that neural networks can approximate continuous operators arbitrarily well—a result generalized by Kovachki et al. (Kovachki2021) for modern neural operator architectures.
@@ 2.2.2 DeepONet: Deep Operator Networks
DeepONet, introduced by Lu et al. (Lu2021) in Nature Machine Intelligence, represents the first practical implementation of operator learning achieving widespread success.
@@
Wang et al. (Wang2021SciAdv) extended DeepONet with physics-informed training, combining data and physical constraints. The loss function:
@@
POD-DeepONet: Bhattacharya et al. (Bhattacharya2021) combined proper orthogonal decomposition with DeepONet. The branch network predicts POD coefficients, and the trunk network reconstructs the solution from basis functions. This reduces output dimensionality and improves data efficiency.
MIONet (Multiple-Input Operator Network): Extends DeepONet to multiple input functions, enabling multi-physics or multi-fidelity (Penwarden2023; Howard2023; Peherstorfer2018) learning (Jin2022).
Fourier-DeepONet: Incorporates Fourier feature embeddings in the trunk network to better capture high-frequency solution components (Li2022).
@@
A breakthrough came from Feng et al. (Nature2025OneShotLearning) in Nature Communications, demonstrating one-shot operator learning from single solution trajectories using self-supervised learning and meta-learning, achieving 5-10% relative errors with just one example—previously requiring hundreds. This is revolutionary for expensive simulations. Yang (Yang2025DeepONet) provided comprehensive generalization bounds for DeepONet with physics-informed training, deriving error bounds in terms of Rademacher complexity. Li et al. (Li2025Architectural) analyzed the trunk-branch architecture and proposed extensions allowing nonlinear interference for improved performance on nonlinear parametric PDEs. Zhong et al. (Zhong2024PICDeepONet) introduced physics-informed compositional DeepONet for complex multi-physics problems with hierarchical decomposition.
@@ 2.2.3 Fourier Neural Operator (FNO)
The Fourier Neural Operator, introduced by Li et al. (Li2021ICLR) at ICLR 2021, represents a breakthrough in operator learning through spectral methods. FNO achieves remarkable computational efficiency and zero-shot super-resolution capabilities by operating in the Fourier domain, making it particularly effective for parametric PDEs on periodic domains or problems amenable to Fourier analysis.
@@
Standard FNO assumes periodic domains and regular grids. Li et al. (Li2023JMLR) extended FNO to irregular geometries through several techniques:
@@
Wang et al. (Wang2022) combined FNO with physics-informed training:
@@
Li et al. (Li2021ICLR), building on earlier work in turbulence modeling (Wang2020Turb; Kurth2023; List2023), demonstrated FNO on 2D turbulent flows governed by the Navier-Stokes vorticity equation:
@@
For inverse problems, Feng et al. (Feng2025SCFNO) introduced Sensitivity-Constrained FNO (SC-FNO) at ICLR 2025, integrating sensitivity analysis into the operator framework for improved parameter inversion. Zhang et al. (Zhang2025PIFNO) developed physics-informed FNO with enhanced constraint enforcement mechanisms.
Conservation laws received special attention: Cardoso-Bihlo and Bihlo (CardoboBihlo2025Conservative) developed exactly conservative physics-informed neural operators in Neural Networks 2025, ensuring discrete conservation of mass, momentum, and energy through learnable adaptive correction. Liu and Tang (Liu2025DiffFNO) combined diffusion models with FNO (DiffFNO) at CVPR 2025 for better uncertainty quantification and robustness, achieving 15-25% accuracy improvements. Kalimuthu et al. (Kalimuthu2025LOGLOFNO) proposed LOGLO-FNO at ICLR 2025 for efficient local-global feature learning. Loeffler et al. (Loeffler2025CMAME) provided rigorous error analysis for FNO on parametric PDEs, establishing convergence rates.
For variable geometries, Zhong and Meidani (Zhong2025PIGANO) developed Physics-Informed Geometry-Aware Neural Operator (PI-GANO) in CMAME 2025, simultaneously generalizing across PDE parameters and domain geometries using signed distance functions and parameter-geometry attention mechanisms, achieving ¡3% relative errors across diverse geometric configurations without expensive FEM data generation.
@@ 2.2.4 Graph Neural Operators and Advanced Architectures
Proposed by Li et al. (Li2020), with extensions in (Li2023GNOT; Hao2023GNOT2; Gao2021; Pfaff2021; Sanchez2020), GNO represents functions on graphs rather than regular grids, naturally handling unstructured meshes arising in finite element simulations or point clouds from experimental measurements.
@@
Graph neural operators have seen significant methodological advances for challenging geometries. Liao et al. (Liao2025Curvature) developed curvature-aware graph attention at ICML 2025 for PDEs on manifolds, incorporating Riemann curvature tensors into message passing for accurate differential equation solving on curved surfaces. Lino et al. (Lino2025Diffusion) introduced diffusion graph networks at ICLR 2025 for complex fluid simulations with irregular boundaries, using learned diffusion processes on graphs. Zou et al. (Zou2025FDIGN) proposed finite-difference-informed graph networks in Physics of Fluids 2025 for incompressible flows on block-structured grids, elegantly combining GNN flexibility with finite-difference structure.
@@
Li et al. (Li2023NeurIPS) introduced MGNO to efficiently handle long-range interactions in graph-based operators. The key innovation is a multipole expansion separating near-field (local) and far-field (global) interactions:
@@
Gupta and Brandstetter (Gupta2022) adapted the U-Net encoder-decoder architecture for operator learning. The multi-scale nature of U-Nets—with skip connections preserving fine details—makes them effective for multi-scale parametric PDEs.
@@
Recent work (Cao2021; Hao2023) explores attention mechanisms for operator learning. The Transolver (Wu2023Transolver; Cao2023) architecture uses:
@@
Emerging paradigms explore fundamentally new approaches to operator learning. Zheng et al. (Zheng2024NIPS) introduced the Mamba Neural Operator at NeurIPS 2024, applying state-space models to PDEs. Mamba uses adaptive state-space matrices with alias-free design for capturing long-range spatiotemporal dependencies at reduced computational cost compared to transformers, achieving competitive performance with better scaling.
Memory-enhanced models address long-time integration challenges. Buitrago-Restrepo et al. (BuitragoRestrepo2025ICLR) demonstrated explicit memory mechanisms for time-dependent PDEs at ICLR 2025. By maintaining memory buffers storing spatiotemporal patterns, they significantly reduced autoregressive errors for long-time integration (500 vs 100 stable time steps, 40% lower error). Le Boudec et al. (LeBoudec2025ICLR) and Moro et al. (Moro2025ICLR) proposed novel attention mechanisms specifically designed for operator learning. Li et al. (Li2025Maxup) introduced max-up training strategies that improve out-of-distribution generalization by augmenting training with worst-case parameter perturbations.

### v3

@@ 1.3.1 Proper Orthogonal Decomposition and Galerkin Projection
The Proper Orthogonal Decomposition (POD), also known as Principal Component Analysis or Karhunen-Loève expansion, forms the cornerstone of traditional ROM approaches (Holmes, 2012; Berkooz et al., 1993). Given snapshot solutions $\{u(\cdot,\cdot;\mu_{i})\}_{i=1}^{N}$ collected at training parameters, POD constructs an optimal low-dimensional approximation subspace.
@@ 1.3.2 Reduced Basis Methods
Reduced basis (RB) methods (Quarteroni et al., 2015; Hesthaven et al., 2016; Rozza et al., 2008) extend POD through sophisticated greedy algorithms for parameter space exploration. The key innovation is adaptive selection of snapshot parameters to maximize approximation quality across $\mathcal{P}$.
@@ 1.3.3 Sparse Grid and Multi-Level Methods
For higher-dimensional parameter spaces, sparse grid methods based on Smolyak’s algorithm (Bungartz and Griebel, 2004) provide a middle ground between full tensor-product grids and random sampling. The key idea is anisotropic refinement, placing more collocation points along directions of high solution variability.
Multi-level Monte Carlo (MLMC) (Giles, 2015) tackles high-dimensional uncertainty quantification by combining low-fidelity (coarse mesh) solutions with high-fidelity corrections, achieving variance reduction without prohibitive cost.
@@ 2.1.1 Foundational Framework
Physics-informed neural networks, pioneered by Raissi et al. (Raissi et al., 2019) in their seminal 2019 work, represent a breakthrough in integrating physical laws with data-driven learning. The core innovation lies in encoding PDE residuals directly into the loss function, enabling networks to discover solutions satisfying both data constraints and governing equations.
@@ 1. Inverse Problems and Parameter Identification
The pioneering work of Raissi et al. (Raissi et al., 2019), building on foundational deep learning approaches for PDEs (Han et al., 2018; Sirignano and Spiliopoulos, 2018; Berg and Nyström, 2018; Yu and others, 2018), demonstrated PINNs’ remarkable capability for inverse problems: inferring unknown parameters from sparse observational data. Complementary methodological developments (Raissi, 2018; Cuomo et al., 2022; Lu et al., 2022a) have further established PINNs as powerful tools for parameter identification. This capability is particularly valuable when parameters represent hidden physical properties.
@@
Case Study: Hidden Fluid Mechanics (Raissi et al., 2020): Raissi et al. demonstrated parameter identification for the Navier-Stokes equations from scattered velocity measurements. Given noisy observations of the velocity field, the method simultaneously reconstructs the full flow field and infers the Reynolds number, achieving relative errors below 1% with measurements at only 0.1% of the domain points.
@@ 2. Adaptive Sampling Strategies
Residual-Based Adaptive Refinement (RAR): At each training iteration, compute residuals at all collocation points and preferentially sample from high-residual regions (Daw et al., 2022; Gao et al., 2021; Lu et al., 2021b; Wu et al., 2020):
@@
Recent work by Wu et al. (Wu et al., 2023a) demonstrated 50-70% reductions in required training data through parameter space active learning (Lookman et al., 2019; Wu et al., 2020) for Reynolds number-dependent flows.
Advanced training methods have improved PINN performance. Wang et al. (Wang et al., 2025) introduced gradient alignment from a second-order optimization perspective, dynamically adjusting loss weights based on gradient alignment to achieve 30-50% training time reductions. De Ryck et al. (De Ryck et al., 2024a) provided an operator preconditioning perspective on physics-informed training. Hwang and Lim (Hwang and Lim, 2024) developed dual cone gradient descent for handling conflicting gradients.
Hao et al. (Hao et al., 2024) released PINNacle at NeurIPS 2024, providing standardized PINN evaluation across 15 canonical PDEs with 20+ method variants. Miyagawa and Yokota (Miyagawa and Yokota, 2024) extended PINN theory to functional differential equations with convergence guarantees.
@@ 3. Multi-Task and Transfer Learning
Desai et al. (Desai et al., 2021), along with comparative studies (Lu et al., 2022a; Geneva and Zabaras, 2022), showed that transfer learning reduces training iterations by up to 80% when adapting PINNs across Reynolds numbers.
@@ 2.1.3 Recent Advances in PINN Architecture and Training
Cao and Zhang (2025) established a fundamental connection between PINN ill-conditioning and the Jacobian matrix condition number of the PDE system. By constructing controlled systems that adjust the condition number while preserving solutions, they demonstrated that reducing the Jacobian condition number leads to faster convergence and higher accuracy. This breakthrough enabled the first successful PINN simulation of three-dimensional flow around the M6 wing at Reynolds number 5,000, representing a significant milestone for industrial-complexity problems. Zhang et al. (2025b) built upon their foundation, applying it to the parametric domain.
@@
To mitigate the curse of dimensionality (CoD) in multi-dimensional PDEs, several separable architectures have been proposed. These methods primarily focus on decomposing the high-dimensional problem. Cho et al. (2023), for instance, introduced Separable Physics-Informed Neural Networks (SPINN), which operates on a per-axis basis to dramatically reduce network forward passes and memory overhead, enabling training with over $10^{7}$ collocation points. Similarly, zhang2025dfs proposed the Dynamic Feature Separation PINN (DFS-PINN), which employs an innovative input-decoupling and dynamic interaction mechanism to achieve significant computational savings. While these methods leverage separation to tackle dimensionality, the concept has also been proven effective against spectral bias. xiong32025separated proposed the Separated-Variable Spectral Neural Network (SV-SNN), which decomposes multivariate functions into products of univariate functions integrated with adaptive Fourier features, specifically to capture high-frequency components effectively.
@@
A paradigm shift in neural architecture comes from Kolmogorov-Arnold Networks (KAN), which replace fixed activation functions with learnable spline-based functions. Wang et al. (2024) proposed Kolmogorov-Arnold-Informed Neural Networks (KINN), systematically comparing KAN and MLP across various PDE formulations including strong form, energy form, and inverse form. KINN demonstrates significant advantages in accuracy and convergence for multi-scale, singularity, stress concentration, and nonlinear hyperelasticity problems, offering better interpretability with fewer parameters.
Building on this foundation, Jacob et al. (2024) introduced Separable Physics-Informed Kolmogorov-Arnold Networks (SPIKANs), applying the separation of variables principle to KANs. By decomposing problems such that each dimension is handled by an individual KAN, SPIKANs drastically reduce computational complexity without sacrificing accuracy, particularly effective for higher-dimensional PDEs where collocation points grow exponentially with dimensionality.
Most recently, Zhang et al. (2025c) proposed Legend-KINN, integrating Legendre orthogonal polynomials into the KAN architecture combined with pseudo-time stepping in backpropagation. Legend-KINN achieves 1–3 orders of magnitude faster convergence than both KAN and MLP under identical parameter settings. Further exploring this direction, xiong22025j developed J-PIKAN, a framework based on Jacobi orthogonal polynomials.
@@ 2.1.4 Advantages and Limitations: A Critical Analysis
4. Flexible Uncertainty Quantification: Bayesian extensions like B-PINNs (Yang et al., 2021) and related uncertainty quantification approaches (Yang and Perdikaris, 2019; Psaros et al., 2023; Lakshminarayanan et al., 2017; Tripathy and Bilionis, 2018) naturally quantify parameter and prediction uncertainties without requiring multiple forward solves.
@@
2. Training Instability: Balancing multiple loss terms in Eq. equation 13 remains challenging. Stiff PDEs, multi-scale problems, and high Reynolds number flows often exhibit training instabilities, with loss components competing rather than cooperating (Krishnapriyan et al., 2021; Wang et al., 2022a).
@@
5. Spectral Bias: Neural networks exhibit spectral bias toward learning low-frequency components of solutions (Rahaman et al., 2019). For parametric problems with high-frequency features (shocks, boundary layers) that vary with parameters, this bias necessitates very deep networks or specialized architectures.
@@ 2.2.1 Mathematical Foundations of Operator Learning
Chen and Chen (Chen and Chen, 1995) established that neural networks can approximate continuous operators arbitrarily well—a result generalized by Kovachki et al. (Kovachki et al., 2023b) for modern neural operator architectures.
@@ 2.2.2 DeepONet: Deep Operator Networks
DeepONet, introduced by Lu et al. (Lu et al., 2021a) in Nature Machine Intelligence, represents the first practical implementation of operator learning achieving widespread success.
@@
Wang et al. (Wang et al., 2021) extended DeepONet with physics-informed training, combining data and physical constraints. The loss function:
@@
POD-DeepONet: Bhattacharya et al. (Bhattacharya et al., 2021) combined proper orthogonal decomposition with DeepONet. The branch network predicts POD coefficients, and the trunk network reconstructs the solution from basis functions. This reduces output dimensionality and improves data efficiency.
MIONet (Multiple-Input Operator Network): Extends DeepONet to multiple input functions, enabling multi-physics or multi-fidelity (Penwarden et al., 2022; Lu et al., 2022b; Peherstorfer et al., 2018) learning (Jin et al., 2022b).
Fourier-DeepONet: Incorporates Fourier feature embeddings in the trunk network to better capture high-frequency solution components (Li et al., 2020c).
@@
A breakthrough came from Jiao et al. (Jiao et al., 2025) in Nature Communications, demonstrating one-shot operator learning from single solution trajectories using self-supervised learning and meta-learning, achieving 5-10% relative errors with just one example—previously requiring hundreds. This is revolutionary for expensive simulations. Yang (Yang, 2025) provided comprehensive generalization bounds for DeepONet with physics-informed training, deriving error bounds in terms of Rademacher complexity. Li et al. (Li et al., 2025a) analyzed the trunk-branch architecture and proposed extensions allowing nonlinear interference for improved performance on nonlinear parametric PDEs. Zhong et al. (Zhong and Meidani, 2024) introduced physics-informed compositional DeepONet for complex multi-physics problems with hierarchical decomposition.
@@ 2.2.3 Fourier Neural Operator (FNO)
The Fourier Neural Operator, introduced by Li et al. (Li et al., 2020a) at ICLR 2021, represents a breakthrough in operator learning through spectral methods. FNO achieves remarkable computational efficiency and zero-shot super-resolution capabilities by operating in the Fourier domain, making it particularly effective for parametric PDEs on periodic domains or problems amenable to Fourier analysis.
@@
Standard FNO assumes periodic domains and regular grids. Li et al. (Li et al., 2023b) extended FNO to irregular geometries through several techniques:
@@
Wang et al. (Wang et al., 2022a) combined FNO with physics-informed training:
@@
Li et al. (Li et al., 2020a), building on earlier work in turbulence modeling (Wang et al., 2017; Kurth et al., 2023; List et al., 2022), demonstrated FNO on 2D turbulent flows governed by the Navier-Stokes vorticity equation:
@@
For inverse problems, Behroozi et al. (Behroozi et al., 2025) introduced Sensitivity-Constrained FNO (SC-FNO) at ICLR 2025, integrating sensitivity analysis into the operator framework for improved parameter inversion. Zhang et al. (Zhang et al., 2025a) developed physics-informed FNO with enhanced constraint enforcement mechanisms.
Conservation laws received special attention: Cardoso-Bihlo and Bihlo (Cardoso-Bihlo and Bihlo, 2025) developed exactly conservative physics-informed neural operators in Neural Networks 2025, ensuring discrete conservation of mass, momentum, and energy through learnable adaptive correction. Liu and Tang (Liu and Tang, 2025) combined diffusion models with FNO (DiffFNO) at CVPR 2025 for better uncertainty quantification and robustness, achieving 15-25% accuracy improvements. Kalimuthu et al. (Kalimuthu et al., 2025) proposed LOGLO-FNO at ICLR 2025 for efficient local-global feature learning. Liu et al. (Liu et al., 2025) provided rigorous error analysis for FNO on parametric PDEs, establishing convergence rates.
For variable geometries, Zhong and Meidani (Zhong and Meidani, 2025) developed Physics-Informed Geometry-Aware Neural Operator (PI-GANO) in CMAME 2025, simultaneously generalizing across PDE parameters and domain geometries using signed distance functions and parameter-geometry attention mechanisms, achieving ¡3% relative errors across diverse geometric configurations without expensive FEM data generation.
@@ 2.2.4 Graph Neural Operators and Advanced Architectures
Proposed by Li et al. (Li et al., 2020b), with extensions in (Li et al., 2023c; Hao et al., 2023; Gao et al., 2022; Pfaff et al., 2021; Sanchez-Gonzalez et al., 2020), GNO represents functions on graphs rather than regular grids, naturally handling unstructured meshes arising in finite element simulations or point clouds from experimental measurements.
@@
Graph neural operators have seen significant methodological advances for challenging geometries. Liao et al. (Liao et al., 2025) developed curvature-aware graph attention at ICML 2025 for PDEs on manifolds, incorporating Riemann curvature tensors into message passing for accurate differential equation solving on curved surfaces. Lino et al. (Valencia et al., 2025) introduced diffusion graph networks at ICLR 2025 for complex fluid simulations with irregular boundaries, using learned diffusion processes on graphs. Zou et al. (Zou et al., 2025) proposed finite-difference-informed graph networks in Physics of Fluids 2025 for incompressible flows on block-structured grids, elegantly combining GNN flexibility with finite-difference structure.
@@
Li et al. (Li et al., 2020d) introduced MGNO to efficiently handle long-range interactions in graph-based operators. The key innovation is a multipole expansion separating near-field (local) and far-field (global) interactions:
@@
Gupta and Brandstetter (Gupta and Brandstetter, 2022) adapted the U-Net encoder-decoder architecture for operator learning. The multi-scale nature of U-Nets—with skip connections preserving fine details—makes them effective for multi-scale parametric PDEs.
@@
Recent work (Cao, 2021a; Hao et al., 2022) explores attention mechanisms for operator learning. The Transolver (Wu et al., 2023b; Cao, 2021b) architecture uses:
@@
Emerging paradigms explore fundamentally new approaches to operator learning. Zheng et al. (Zheng et al., 2024) introduced the Mamba Neural Operator at NeurIPS 2024, applying state-space models to PDEs. Mamba uses adaptive state-space matrices with alias-free design for capturing long-range spatiotemporal dependencies at reduced computational cost compared to transformers, achieving competitive performance with better scaling.
Memory-enhanced models address long-time integration challenges. Buitrago-Restrepo et al. (Buitrago et al., 2025) demonstrated explicit memory mechanisms for time-dependent PDEs at ICLR 2025. By maintaining memory buffers storing spatiotemporal patterns, they significantly reduced autoregressive errors for long-time integration (500 vs 100 stable time steps, 40% lower error). Le Boudec et al. (Le Boudec et al., 2025) and Moro et al. (Moro and Chamon, 2025) proposed novel attention mechanisms specifically designed for operator learning. Li et al. (Li et al., 2025b) introduced max-up training strategies that improve out-of-distribution generalization by augmenting training with worst-case parameter perturbations.